# Advancing Neural Network Performance through Emergence-Promoting Initialization Scheme

### v1

@@ 3 Method
On learning rates
@@
Empirically, the scaling factor $\alpha$ is also dependent on learning rate. When the learning rate is faster, more stability is usually required and hence we need smaller emergence. There has been some theoretical results on the optimal learning rate, for example, Hettinger et al. (2020) suggested that the optimal learning rate should be inversely proportional to the gradient magnitude at initialization:
@@
For a two layer network, if we choose $\alpha_{0}=2$ for learning rate $\eta_{0}=1e-3$, then for a different learning rate $\eta=1e-4$ we should choose $\alpha=6.32$.
@@
Other architectures and block wise initialization
Note that the mathematical equation (2) only works for MLP. For convolutional layers, we have the following:
@@
In most convolutional architectures and transformers, there are MLP blocks presented. While we can only do our intilization to the MLP blocks and see an improvement, we can also consider applying scheme (8) to the convolutional layers. Given this formula for emergence, we can increase the global emergence, but also increase the local emergence by applying scheme (8) to some of the layer blocks. We will discuss this in the next section.
4 experiments
We evaluate our initialization scheme on benchmark datasets for image classification and machine translation tasks. For image classification, five different architectures are evaluated for CIFAR10 [37]. For machine translation, we use our initialization scheme to find good initializations for a Post-LN Transformer without any change to its original architecture on IWSLT-14 De-En [39].
We conduct our experiments in PyTorch. We use the fairseq library for machine translation [40]. All the experiments on CIFAR-10 and IWSLT-14 DE-EN can run with one single NVIDIA A100 GPU with 13GB of RAM. Our initialization scheme first initializes the weights using Kaiming initialization for all the Conv and MLP layers for image classification. For machine translation, we use the default Xavier initialization [12]. Base on the discussion in the previous section, we choose the scale factors $\alpha=2$ with out batch normalization and $\alpha_{i}=5$ with batch normalization.
On CIFAR-10, we focus on MLP and the feedforward VGG net with and without BN layers. Since ResNet has recursive network structure, we leave it to a future work to establish the emergence formula on it. For MLP, we use a simple MLP architecture with $3$ hidden layers. For VGG net, we use VGG-19 and our initialization scheme is compared with four different methods/settings: 1) Kaiming Initialization [13]; 2) First train the network for one epoch with a constant learning rate equal to the starting learning rate, labelled as “+1 epoch (Const. LR)" in Table 3; 3) First train the network for one epoch with a linear warmup learning rate, labbeled as “+1 epoch (Warmup)" in Table 3; 4) MetaInit [27]. The data is from GradInit[8]. On CIFAR-10, we train networks with a batch size of 128, and a constant learning rate of $0.001$.
From our experiments we can see BN does stabilize VGG-19 and allows training with stronger emergence (larger value of $\alpha$).
(a) w/o batch normalization,$\alpha=2$
(b) w/ batch normazalization, $\alpha=10$
Figure 3: Training loss and test accuracy of MLP on CIFAR-10.
Model VGG-19 VGG-19 ResNet-110 ResNet-110
w/o BN w/ BN w/o BN w/ BN
Kaiming $Acc_{1}$ $29.1\pm 1.5$ $12.6\pm 0.6$ $16.1\pm 2.1$ $23.2\pm 0.9$
+1 epoch (Const. LR) $Acc_{1}$ $37.2\pm 1.1$ $19.6\pm 4.0$ $21.0\pm 3.8$ $32.5\pm 3.8$
+1 epoch (Warmup) $Acc_{1}$ $37.4\pm 1.2$ $53.5\pm 2.9$ $19.8\pm 0.5$ $48.7\pm 1.1$
MetaInit $Acc_{1}$ $30.5\pm 0.9$ $35.1\pm 0.6$ $14.6\pm 2.2$ $29.0\pm 1.5$
GradInit $Acc_{1}$ $29.3\pm 0.6$ $47.8\pm 1.8$ $36.2\pm 0.8$ $38.2\pm 0.9$
Ours $Acc_{1}$ $46.2\pm 0.6$ $52.4\pm 1.0$ $45.3\pm 2.0$ $48.0\pm 1.5$
Table 2: Evaluate the transformer model for 80 epochs
Method $\text{BLEU}_{1}$ $\text{BLEU}_{best}$
Xavier - 34.85
T-Fixup 3.96 34.78
Ours 4.8 35.13
We can see the since batch normalization promotes a good stability, we are free to choose larger $\alpha$.
IWSLT’14 DE-EN [28] is a German to English translation dataset that has 160k training examples. Our Transformer model is inherited from [3], which is a Post-LN Transformer placing its Layer Normalization after the summation of the skip connection and the residual branch. It has a $512$- dimensional word embedding layer and 1024 dimensions in its hidden FFN layer. It has $6$ encoder layers and $6$ decoder layers. We choose the learning rate to be $45e-$ 4 with inverse-sqrt learning schedule with $4000$ warmup updates and weight decay of $0.0001$.
Based on the discussion in the previous section, we have two ways of promoting emergence: A) promote emergence globally, by reducing the magnitude of weights in the encoder layers and increasing the magnitude of weights in the decoder layers; B) promote emergence clockwise, where we apply equation (8) to each encoder/ decoder block. In our transformer architecture, there is MLP block in each encoder/decoder layer consisting of $2$ layers. In our experiments, we can choose $\alpha$ up to $10$ for each MLP block, and notably we can see fast increase of BLEU score in the first few epochs. For example, with $\alpha=10$, the BLEU score after first epoch reaches $6.02$ while for T-Fixup the BLEU score after first epoch is $3.79$.
5 Conclusion
In this paper, we introduced a novel and straightforward neural network initialization scheme inspired by the concept of emergence. Building on the emergent network measures proposed by Li (2023) [2], our method adjusts the layer-wise variance parameters to enhance the number of paths from inactive to active nodes, thereby achieving higher emergence values. This approach is not only easy to implement but also requires no additional optimization or training steps compared to conventional methods like Xavier and Kaiming initialization.
Our extensive evaluations across various architectures, including convolutional neural networks (CNNs) for image recognition and transformers for machine translation, demonstrate the significant advantages of our initialization scheme. The empirical results show that our method substantially improves model accuracy and convergence speed on standard datasets such as CIFAR-10 and the IWSLT-14 translation task.
By focusing on enhancing the emergent properties of neural networks from the outset, our initialization scheme facilitates better feature differentiation and integration, leading to more effective learning. The theoretical foundation and practical benefits of our method offer a promising direction for future research and applications in machine learning.
Our work contributes to the growing body of research that seeks to understand and harness the power of emergence in neural networks. By providing a simple yet powerful modification to existing initialization techniques, we open new avenues for improving neural network training methodologies. Future work could explore further optimizations and adaptations of our initialization scheme to other types of neural architectures and more complex tasks.
Overall, our emergence-promoting initialization scheme represents a valuable addition to current neural network initialization practices, offering both theoretical insights and practical improvements for the development of more robust and capable machine learning models.

### v3

@@ 3 Method
4 Experiments
We evaluate our initialization scheme on benchmark datasets for image classification and machine translation tasks. For image classification, five different architectures are evaluated for CIFAR-10 (Krizhevsky et al., 2009), and ResNet-50 (He et al., 2016) is evaluated for ImageNet(Deng et al., 2009). For machine translation, we use our initialization scheme to find good initializations for a Post-LN Transformer without any change to its original architecture on IWSLT-14 De-En (Cettolo et al., 2014).
We conduct our experiments in PyTorch. We use the fairseq library for machine translation (Ott, 2019). All the experiments on CIFAR-10 and IWSLT-14 DE-EN can run with one single NVIDIA A100 GPU. Our initialization scheme first initializes the weights using Kaiming initialization for all the Conv and MLP layers for image classification. On ImageNet, we compare with Kaiming Initialization, and GratInit. Each block in ResNet-50 is initialized independently with $\alpha=2$. We use batch normalization to increase stability.
For machine translation, we use the default Xavier initialization (Glorot & Bengio, 2010). Base on the discussion in the previous section, we choose the scale factors $\alpha=2$ with out batch normalization and $\alpha_{i}=5$ with batch normalization.
On CIFAR-10, we focus on MLP and the feedforward VGG net with and without BN layers. Since ResNet has recursive network structure, we leave it to a future work to establish the emergence formula on it. For MLP, we use a simple MLP architecture with $3$ hidden layers. For VGG net, we use VGG-19 and our initialization scheme is compared with four different methods/settings: 1) Kaiming Initialization (He et al., 2015); 2) First train the network for one epoch with a constant learning rate equal to the starting learning rate, labelled as “+1 epoch (Const. LR)" in Table 1; 3) First train the network for one epoch with a linear warmup learning rate, labbeled as “+1 epoch (Warmup)" in Table 1; 4) MetaInit (Dauphin & Schoenholz, 2019). The data is from GradInit(Zhu et al., 2021). On CIFAR-10, we train networks with a batch size of 128, and in our initialization scheme, we adopt a constant learning rate of $0.001$, while in other initialization models, much larger learning rate (for example, $0.1$) has been adopted. Our scheme has significant performance even though the learning rate is much smaller.
From our experiments we can also see BN does stabilize VGG-19 and allows training with stronger emergence (larger value of $\alpha$). This shows the particular promising application of our scheme combined with batch normalzation. We can see from our numerical simulation that since batch normalization promotes good stability, we are free to choose larger $\alpha$.
(a) w/o batch normalization,$\alpha=2$
(b) w/ batch normazalization, $\alpha=10$
Figure 3: Training loss and test accuracy of MLP on CIFAR-10.
Model VGG-19 VGG-19 ResNet-110 ResNet-110 ResNet-1202
w/o BN w/ BN w/o BN w/ BN w/ BN
Kaiming $29.1\pm 1.5$ $12.6\pm 0.6$ $16.1\pm 2.1$ $23.2\pm 0.9$ $12.9\pm 2.8$
+1 epoch (Const. LR) $37.2\pm 1.1$ $19.6\pm 4.0$ $21.0\pm 3.8$ $32.5\pm 3.8$ $12.6\pm 2.8$
+1 epoch (Warmup) $37.4\pm 1.2$ $53.5\pm 2.9$ $19.8\pm 0.5$ $48.7\pm 1.1$ $28.1\pm 1.3$
MetaInit $30.5\pm 0.9$ $35.1\pm 0.6$ $14.6\pm 2.2$ $29.0\pm 1.5$ $11.7\pm 1.6$
GradInit $29.3\pm 0.6$ $47.8\pm 1.8$ $36.2\pm 0.8$ $38.2\pm 0.9$ $29.0\pm 1.1$
Ours $\textbf{46.2}\pm 0.6$ $52.4\pm 1.0$ $\textbf{45.3}\pm 2.0$ $48.0\pm 1.5$ $\textbf{29.8}\pm 1.7$
Table 2: Accuracy after epoch 1 of ResNet-50 models on ImageNet. Results from (Zhu et al., 2021).
Model Kaiming GradInit Ours
$Acc_{1}$ 14.6 19.2 23.2
Table 3: A comparison of Emergence-Promoting Initialization with other initialization for training the Post-LN Transformer model on the IWSLT-14 De-EN dataset. (Evaluate after 80 epochs)
Model $\text{BLEU}_{1}$ $\text{BLEU}_{best}$
Xavier - 34.85
T-Fixup 3.96 34.78
Ours 4.8 35.13
IWSLT’14 DE-EN (Cettolo et al., 2014) is a German to English translation dataset that has 160k training examples. Our Transformer model is inherited from (Vaswani, 2017), which is a Post-LN Transformer placing its Layer Normalization after the summation of the skip connection and the residual branch. It has a $512$- dimensional word embedding layer and 1024 dimensions in its hidden FFN layer. It has $6$ encoder layers and $6$ decoder layers. We choose the learning rate to be $45e-$ 4 with inverse-sqrt learning schedule with $4000$ warmup updates and weight decay of $0.0001$.
Based on the discussion in the previous section, we have two ways of promoting emergence: A) promote emergence globally, by reducing the magnitude of weights in the encoder layers and increasing the magnitude of weights in the decoder layers; B) promote emergence clockwise, where we apply (8) to each encoder/ decoder block. In our transformer architecture, there is MLP block in each encoder/decoder layer consisting of $2$ layers. In our experiments, we can choose $\alpha$ up to $10$ for each MLP block, and notably we can see fast increase of BLEU score in the first few epochs. For example, with $\alpha=10$, the BLEU score after first epoch reaches $6.02$ while for T-Fixup the BLEU score after first epoch is $3.79$.
In all our numerical experiments, we notice our initialization leads to better training performance when combined with batch normalization, weight decay and other techniques that promotes stability and prevents over-fitting, while in other cases the model could be trapped in local minimum. We encourage researchers to combine our initialization scheme with other stability-promoting considerations, which could potentially further improve the performance (especially long term) of our initialization.
5 Conclusion
In this paper, we introduced a novel and straightforward neural network initialization scheme inspired by the concept of emergence. Building on the emergent network measures proposed by (Li et al., 2023), our method adjusts the layer-wise variance parameters to enhance the number of paths from inactive to active nodes, thereby achieving higher emergence values. This approach is not only easy to implement but also requires no additional optimization or training steps compared to conventional methods like Xavier and Kaiming initialization. Our extensive evaluations across various architectures, including convolutional neural networks (CNNs) for image recognition and transformers for machine translation, demonstrate the significant advantages of our initialization scheme. The empirical results show that our method substantially improves model accuracy and convergence speed on standard datasets such as CIFAR-10, ImageNet and the IWSLT-14 translation task.
Our work contributes to the growing body of research that seeks to understand and harness the power of emergence in neural networks. By providing a simple yet powerful modification to existing initialization techniques, we open new avenues for improving neural network training methodologies. Future work could explore further optimizations and adaptations of our initialization scheme to other types of neural architectures and more complex tasks. Overall, emergence-promoting initialization scheme represents an addition to current neural network initialization practices, offering both theoretical insights and practical improvements for the development of more robust and capable machine learning models.
Appendix A Appendix
A.1 Mathematical Representation of Neural Networks
To perform mathematical computation of emergence, we use quiver representation as the representation of a neural network. Formally, a quiver is a directed graph where loops and multiple arrows between two vertices are allowed, defined as follows:
  • A quiver $Q$ is a quadruple $Q=(Q_{0},Q_{1},h,t)$ where $Q_{0}$ is a finite set of vertices, $Q_{1}$ is a finite set of arrows, and $h$ and $t$ are functions $Q_{1}\rightarrow Q_{0}$. For an arrow $a\in Q_{1}$, $h(a)$ and $t(a)$ are called the head and tail of $a$.
  • We get a representation $V$ of $Q=(Q_{0},Q_{1},h,t)$ if we attach to every vertex $x\in Q_{0}$ a finite dimentional vector space $V(x)$ and to every arrow $a\in Q_{1}$ a linear map $V(a):V(ta)\rightarrow V(ha)$.
Quiver representation can be used to model the dynamics on the network Derksen & Weyman (2017); Armenta & Jodoin (2021); Armenta et al. (2023). We provide two examples of quiver representation in Figure Figure 4.
Figure 4: Additional examples of quivers. (a): A quiver $Q$ with vertices $V=\{a,b,c,d\}$ and oriented edges $E=\{\alpha,\beta,\gamma,\delta\}$, (b) and (c): two quiver representations over $Q$. Adapted from (Derksen & Weyman, 2017).
Theorem (Proposition 5.3 in (Li et al., 2023)) Given the functor $\Phi$ which preserves partial structure in a quiver representation $W$ by deleting a set of edges $E$, the derived functor of $\Phi$ is
$R^{1}\Phi(W)=\bigoplus_{a\in E}\Phi(W(ta)\otimes P_{ha})$ (11)
where $ta$ is the tail of edge $a$ (the starting node), $ha$ is the head of edge $a$ (the ending node), $W(ta)$ is the vector space associated to node $ta$, $P_{ha}$ is the vector space spanned by all paths originating from node $ha$.
Proof.
Based on (Derksen & Weyman, 2017), for representation $W$ in $\textbf{Rep}(Q)$ we have the projective resolution
00{0}${W}$${\bigoplus\limits_{x\in Q_{0}}W(x)\otimes P_{x}}$${\bigoplus\limits_{a\in Q_{1}}W(ta)\otimes P_{ha}}$00{0}$\scriptstyle{f^{W}}$$\scriptstyle{d^{W}}$ (12)
where
$f^{W}:\bigoplus\limits_{x\in Q_{0}}W(x)\otimes P_{x}\rightarrow W$ (13)
is defined by
$f^{W}(w\otimes p)=p\cdot w,$ (14)
and
$d^{W}:\bigoplus\limits_{a\in Q_{1}}W(ta)\otimes P_{ha}\rightarrow\bigoplus\limits_{x\in Q_{0}}W(x)\otimes P_{x}$ (15)
is defined by
$d^{W}(w\otimes p)=(a\cdot w)\otimes p-w\otimes pa.$ (16)
Now we compute the first left derived functor $R^{1}\Phi$. By definition (Rotman & Rotman, 2009), it is the 1st homology object of the sequence above under the image $\Phi$, formally, $R^{1}\Phi=\ker\Phi d^{W}$, where $d^{W}$ is defined in (15). Now if an edge $a$ is deleted by the functor $\Phi$ then for any $w\in W(ta)$ and $p\in P_{ha}$, we have$(a\cdot w)\otimes p=w\otimes ap=0$, hence $\Phi(W(ta)\otimes P_{ha})\subseteq\ker\Phi d^{W}$. If $a$ is preserved under $\Phi$, then $\Phi d^{W}$ will act the same as $d^{W}$ on $\Phi(W(ta)\otimes P_{ha})$, and $d^{W}$ is injective due to the exactness of resolution, $\Phi(W(ta)\otimes P_{ha})$ will be non-zero thus not contribute to $\ker\Phi d^{W}$.
This theorem computes $R^{1}\Phi$, which evaluates the difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$, thus encodes the potential of a system for emergence. Taking advantage of this theorem, we can take the dimension of $R^{1}\Phi(W)$ as a numerical approximation of the potential for emergence of $W$ when the network interact with other networks:
$\dim R^{1}\Phi_{l}(W)$ $=\dim\bigoplus_{e\in E}\Phi_{l}(W(he)\otimes I_{te})$
$=\sum_{e\in E}\dim\Phi_{l}(W(he))\times\dim\Phi_{l}(I_{te}).$ (17)
Here $\dim\Phi_{r}(V(te))$ and $\dim\Phi_{l}(W(he))$ is the dimension of the image of the vector space $V(te)$ and $W(he)$ under the functor, and $\dim\Phi_{r}(P_{he})$ and $\dim\Phi_{l}(I_{te})$ is the dimension of the image of the path algebra $P_{he}$ and $I_{te}$ under the functor.
Given a network $G$, and a sub-network $H$ which represents its effect or observation under $\Phi$, where their relation are shown as follows:
[…]
then we have the following measure of emergence for networks:
$\text{Emergence}(G,H)$ $=\sum_{x\in G\setminus H}\#\text{paths in $H$ from $N_{H}(x)$ to $H$},$ (18)
where $H$ represents the part of network structure being preserved by $\Phi$, the partial observation. $N_{H}(x)$ is the set of downstream neighbors of $x$ in $H$.
A.2 Hyperparameters
@@
On learning rates Empirically, the scaling factor $\alpha$ is also dependent on learning rate. When the learning rate is faster, more stability is usually required and hence we need smaller emergence. There has been some theoretical results on the optimal learning rate, for example, (Hettinger, 2019) suggested that the optimal learning rate should be inversely proportional to the gradient magnitude at initialization:
@@
For a two layer network, if we choose $\alpha_{0}=2$ for learning rate $\eta_{0}=0.001$, then for a different learning rate $\eta=0.0001$ we should choose $\alpha=6.32$.
@@
Other architectures and block wise initialization Note that (9) only works for MLP, which has good symmetry. For convolutional layers, we can modify (9) to get the following measure of emergence:
@@
In most convolutional architectures and transformers, there are MLP blocks presented. While we can only do our internalization to the MLP blocks and see an improvement, we can also consider applying scheme (19) to the convolutional layers. Given this formula for emergence, we can increase the global emergence, but also increase the local emergence by applying scheme (19) to some of the layer blocks. We will discuss this in the next section.